OpenStudy (anonymous):

Solve. 4over7=21overx-3 i need help with this

OpenStudy (♪chibiterasu):

\[4/7 = 21/x-3 \]

OpenStudy (anonymous):

yea just like that but i need help

OpenStudy (♪chibiterasu):

Well, we cross mulitply

OpenStudy (♪chibiterasu):

Do you know how to do that?

OpenStudy (anonymous):

can u help me

OpenStudy (♪chibiterasu):

Cross multiply: 4(x - 3) = 4x - 12 7•21 = 147

OpenStudy (♪chibiterasu):

http://www.youtube.com/watch?v=Uv_hPQfJE4g

OpenStudy (♪chibiterasu):

This should help you :3

OpenStudy (♪chibiterasu):

4x - 12 = 147

OpenStudy (♪chibiterasu):

Now we can isolate x.

OpenStudy (♪chibiterasu):

First add 12 from both sides to remove the 12

OpenStudy (♪chibiterasu):

4x - 12 + 12 = 147 + 12 4x = 159 Now divide 4 from bot sides: 4x/4 = 159/4 x = 39.75

OpenStudy (♪chibiterasu):

It's solved now . . .

OpenStudy (♪chibiterasu):

Your welcome . . .

OpenStudy (♪chibiterasu):

Lol nvm x3

OpenStudy (anonymous):

thank u